# Filtrated Grouping in Multiple Functional Regression

### v2

@@ Abstract
To understand and communicate the risk of chronic joint disease associated with aging, it is essential to investigate how age is associated with gait patterns, particularly through gait angular kinematics. Motivated by this need, and by the critical role of joint coordination in gait, we propose a novel covariate grouping framework within the context of multiple functional regression, where a scalar response is linked to multiple functional covariates. We apply this approach to investigate the relationship between chronological age and gait angular kinematics, aiming to uncover biomechanical patterns that signal age-related gait pattern evolution. Specifically, we develop a forest-structured covariate grouping framework in which different functional covariates are aggregated hierarchically based on the level of coefficient homogeneity. This approach allows for the analysis of both common and idiosyncratic effects of covariates in a nuanced, multi-resolution manner. The identification of the forest structure is entirely data-driven and requires no prior knowledge, providing valuable insights into the interdependence among covariates. Compared to existing methods, the proposed regression framework demonstrates superior predictive power and offers more insightful interpretability on joint coordination. In addition, the proposed framework is broadly applicable and can be readily extended to analyze multivariate functional data in other scientific domains.
@@ 1 Introduction
1.1 Backgrounds and Objectives
@@
Gait analysis (GA) is a widely utilized method for examining and understanding angular kinematics, which involves studying the rotational movements of joints during activities such as walking. This analysis provides valuable insights into joint function and coordination, aiding in the assessment of movement patterns and the identification of potential abnormalities (see e.g., Kay et al. [21], Lofterød et al. [26] and Rao et al. [33]). The study in this paper is driven by a growing scientific interest in the intersection of chronic risk communication and gait analysis.
@@
As individuals age, changes in joint flexibility, muscle strength, and neuromuscular control can alter these angular movements, leading to decreased mobility and increased fall risk. An emerging concept, referred to as “effective age” has garnered increasing attention in recent years (see e.g., Anderson and Loeser [2], Loeser [25], Li et al. [24], and Spiegelhalter [37]). Effective age represents the age of a “healthy” individual who shares the same risk profile as a person being evaluated. In many cases, a person’s health status is more accurately reflected by their effective age rather than their chronological age, as it accounts for the impact of chronic disease risks and overall well-being. Evaluating effective age through gait analysis is a vital component in the proactive management of joint health, facilitating early detection of degeneration, customization of rehabilitation strategies, and the promotion of sustained mobility and independence.
@@
Here we analyze a dataset generated from the experiment conducted by Fukuchi et al. [15]. The study involves 42 healthy volunteers aged between 21 and 84, all of whom had no lower-extremity injuries in the six months prior to participation and no orthopedic or neurological conditions affecting their gait. Studying the association between chronological age and gait angular kinematics in healthy individuals is essential, as it serves as a baseline for assessing new subjects. Data were collected using standard gait-analysis procedures in a $10\mbox{m}\times 12\mbox{m}$ room equipped with 12 cameras, five force platforms embedded in the floor, and a dual-belt instrumented treadmill. This study focuses on treadmill walking data, as they offer greater consistency and are not influenced by terrain variations. The angular kinematics were recorded at various gait speeds, categorized into eight groups (S1 — S8) ranging from the slowest to the fastest speeds. The ensemble averages of angular kinematics across different speed levels are shown in Figure 1. We develop an innovative tool to study the association between joint angular kinematics and chronological age, as a building block for effective age evaluation. To achieve this, we employ multiple functional regression, where age serves as the response variable and angular kinematics curves are the predictors.
@@
Figure 1: Angular kinematics during treadmill walking averaged over different levels of gait speed (right-hand side). S5 represents the comfortable walking speed.
@@ 1.2 Functional Linear Regression and Covariate Grouping
Multivariate functional data analysis (FDA) offers a powerful toolbox for handling multiple complex functional objects. Its primary strength lies in addressing the dual challenges of high dimensionality inherent in individual functional objects and the intricate interdependence among them. Recent advancements in this field have enabled researchers to tackle diverse challenges across various domains, including neuroimaging analysis (e.g., Wang et al. [38] and Jiao et al. [19]), public health (e.g., Zhou et al. [42]), and gait analysis (e.g., Cheng et al. [4], Dannenmaier et al. [8], and Cao et al. [3]). These developments underscore the versatility and potential of multivariate FDA in addressing complex, high-dimensional problems across diverse domains. Among its various tools, functional linear regression (see Ramsay and Silverman [32] and Morris [28] for overviews) stands out as a well-established class of linear models. In this article, we focus on the following model with a scalar response $y_{n}\in\mathbb{R}$ and multiple functional covariates $\{X_{jn}(t)\in L^{2}[0,1]\colon j=1,\ldots,p\}$,
@@
In gait analysis, coordination among different joints is pivotal for understanding human movement patterns and their underlying biomechanics. Existing research underscores the interdependence of angular kinematics, where the motion of one joint often influences or compensates for the movement of others to maintain balance, stability, and forward progression (see e.g., Ren et al. [34] and Dussault-Picard et al. [10]). Thus, an important task is to account for the interdependence among different covariates (i.e., the angular kinematics of various joints). Such interdependence often leads to covariates being associated with the response variable (e.g., chronological age) in a homogeneous manner. A reasonable strategy, therefore, is to group covariates based on the similarity of their effects on the response. Notable works in this area include Shen and Huang [36], Ke et al. [22], Ma and Huang [27], She et al. [35], and Jiao and Chan [18]. However, ordinary grouping pursuit methods face a fundamental limitation: they rely on a uniform grouping structure. For functional data, this restriction is inappropriate due to its infinite-dimensional nature. Specifically, different components of functional covariates may exhibit distinct grouping structures. Consequently, assuming a common structure can lead to model misspecification, reduced estimation efficiency, and hindered interpretability. Therefore, adopting flexible grouping strategies that accommodate variations in grouping structures across different dimensions is essential for accurate modeling and insightful interpretation.
@@ 1.3 Contributions and Novelties
The primary methodological contribution of this paper is the development of a filtration grouping algorithm for model (1-1), termed filtrated grouping, which allows for varying grouping structures across different dimensions. Here, we define “homogeneous components” as covariate components that share a common coefficient score in some dimension, reflecting their collective contribution to the response variable. The key idea is to iteratively filter out the homogeneous components of functional covariates that exhibit significant associations with the response variable, thereby constructing dimension-specific grouping structures. In the application, we employ the proposed methodology to examine the relationship between angular kinematics and chronological age for the healthy participants, with a particular focus on the coordination among different joints. The multi-resolutional coordination structure is discovered and highlights several important factors in the evaluation of effective age. The findings contribute to a deeper understanding of angular kinematics and offer a valuable tool for evaluating joint effective age, providing insights for enhancing age-related gait assessments and joint health evaluations.
This idea is inspired by filtration, a widely used technique in graph analysis and topological data analysis (see, e.g., Edelsbrunner et al. [11], Lee et al. [23], and Hofer et al. [16]), which iteratively constructs a sequence of nested structures and serves as a fundamental building block for multi-resolution analysis. However, it has been rarely applied in multivariate functional data analysis. To the best of our knowledge, the only related work is Jiao et al. [19], which employs filtration to construct a multi-resolution structure for partial common functional principal components (PCfPC), offering a more flexible and efficient representation of PCfPCs. This paper is the first work to develop filtration techniques for multiple functional regression.
The rest of the article is organized as follows. We present the filtrated grouping structure, and develop the model estimation procedure in Section 2. The identification of the filtrated grouping structure is discussed in Section 3. Simulation results are presented in Section 4. In Section 5, we apply our methods to analyze the association between chronological age and angular kinematics data. The paper is concluded in Section 6.
2 Filtrated multiple functional regression
@@ 2.1 Preliminaries
[…]
@@
in which the basis functions $\{\nu_{jd}(t)\colon d\geq 1\}$ and $\{u_{jd}(t)\colon d\geq 1\}$ can either be orthonormal or not. The term $M_{j}=\{M_{j,dd^{\prime}}\colon d,d^{\prime}\geq 1\}$ represents an infinite matrix satisfying the condition $\sum_{d^{\prime}\geq 1}M_{j,d_{1}d^{\prime}}(\int\nu_{jd_{2}}u_{jd^{\prime}})=\bm{1}_{\{d_{1}=d_{2}\}}$. Then model (1-1) can be equivalently rewritten in the following form
@@
For illustrating the filtrated grouping structure, we focus on the equivalent form (2-2), as the representation is dimension-specific (layer-specific), enabling a more precise characterization of grouping structures across different dimensions.
@@ 2.2 Filtrated Grouping Structure
Given a grouping structure $G=(\mathcal{K}_{1},\ldots,\mathcal{K}_{m})$ satisfying $\mathcal{K}_{i}\cap\mathcal{K}_{i^{\prime}}=\emptyset$ for $i\neq i^{\prime}$ and $\cup_{i=1}^{m}\mathcal{K}_{i}=\{1,\ldots,p\}$, the ordinary grouped model based on coefficient homogeneity is
@@
where $Z_{in}(t)=\sum_{j\in\mathcal{K}_{i}}X_{jn}(t)$ and $\alpha_{i}(t)$ is the common coefficient function of group $\mathcal{K}_{i}$.
The ordinary model (1-1) cannot account for the homogeneity of functional coefficients and the estimation burden is high even for moderate $p$ (see e.g., Jiao and Chan [18]). Although the grouped model (2-3) alleviates the estimation burden and improve model interpretability, it fails to capture variations in grouping structures across different dimensions. These limitations motivate the development of a more flexible modeling framework that can accommodate dimension-specific grouping variability.
Building upon the form (2-2), we develop the following filtrated grouping structure
@@
where $\mathcal{K}_{d,i}$ denotes the $i$-th group in the $d$-th dimension/layer and $m_{d}$ is the number of groups in the $d$-th dimension. We define the homogeneous components as follows
$\zeta_{nd}^{(i)}\overset{\vartriangle}{=}\sum_{j\in\mathcal{K}_{d,i}}\xi_{jn,d},\ i\geq 1,\ d\geq 1,$
and aim to estimate the associated coefficient scores $\{a_{d,i}\colon i\geq 1,\ d\geq 1\}$.
The coefficient homogeneity is captured in a multi-resolution manner, resulting in group indices $\{\mathcal{K}_{d,i}\colon i\geq 1,d\geq 1\}$ that follow a forest structure (see Figure 2). Compared to the ordinary grouped model (2-3), the filtrated grouped model (2-4) accommodates varying grouping structures across different dimensions. In the first layer, we identify the most homogeneous covariate components and group them accordingly to obtain $\{\zeta_{n1}^{(i)}\colon i\geq 1\}$ and estimate $\{a_{1,i}\colon i\geq 1\}$. After filtering out these homogeneous components from each covariate, we divide $\{\mathcal{K}_{1,i}\colon i\geq 1\}$ into $\{\mathcal{K}_{2,i}\colon i\geq 1\}$, which are either nested within or identical to the groups in the previous layer (i.e., for any $i\geq 1$, $\mathcal{K}_{2,i}\subseteq\mathcal{K}_{1,i^{\prime}}$ for some $i^{\prime}\geq 1$), and then obtain the homogeneous components in the new layer (i.e., $\{\zeta_{n2}^{(i)}\colon i\geq 1\}$) and estimate the associated coefficient scores (i.e., $\{a_{2,i}\colon i\geq 1\}$). The splitting procedure continues until a predefined stopping criterion is met. In the deeper layers of the filtration, higher-resolution homogeneous components, constructed from fewer covariates, are progressively extracted.
Figure 2: Hierarchical forest structure of coefficient homogeneity. Here, $L_{d}$ signifies the $d$-th layer of filtration. Different groups are represented by different colored solid circles. Each group is either nested within or identical to its “parent” group in the previous layer, i.e., the group from which it originated through a split.
In the remainder of the article, we address the following problems: How to determine a reasonable forest structure for covariate grouping? Given the forest structure, how to identify homogeneous components and estimate coefficient scores? In Section 3, we construct the forest structure based on the grouping path derived from the grouping pursuit problem, and in Section 2.3, we address the second problem by developing a filtrated functional partial least squares approach.

### v3

# Filtrated Kinematic Connectivity Analysis for Lower‑limb Joint Effective Age Evaluation

@@ Abstract
To understand and communicate the risk of chronic lower-limb joint diseases associated with aging, it is crucial to investigate the relationship between age and gait dynamics, particularly through angular kinematics. One key challenge is that angular kinematic trajectories are highly interconnected, and the structures of the interconnections vary across different components. Neglecting the interconnections and the variability in the connectivity structures impairs the understanding of age‑associated gait coordination. To this end, we develop a novel kinematic connectivity analysis framework, grounded in multiple functional regression, to evaluate lower-limb joint effective age and uncover age-related kinematic features. The proposed approach is built upon the concept of filtration, a widely used tool in network analysis and topological data analysis for multi-resolution exploration. Specifically, we develop a forest-structured covariate grouping framework in which different kinematic trajectories are aggregated hierarchically to capture both (partially) shared and idiosyncratic motion signatures which are strongly associated with aging. We also develop a novel filtrated functional partial least squares approach for model estimation and feature extraction. Compared to existing approaches, our proposed approach demonstrates superior predictive power while providing novel insights into the coordinated evolution of angular kinematics during aging. In addition, the proposed framework is broadly applicable and can be readily extended in other scientific domains.
@@ 1 Introduction
1.1 Backgrounds, Data Description, and Objectives
@@
Gait analysis (GA) is a widely utilized method for examining and understanding angular kinematics, which involves studying the rotational movements of lower-limb joints during activities such as walking. This analysis provides valuable insights into joint function and coordination, aiding in the assessment of movement patterns and the identification of potential abnormalities (see e.g., Kay et al. [20], Lofterød et al. [25] and Rao et al. [31]). As individuals age, changes in joint flexibility, muscle strength, and neuromuscular control can alter these angular movements, leading to decreased mobility and increased fall risk. The study in this paper is driven by a growing scientific interest in the intersection of chronic risk communication and gait analysis.
@@
An emerging concept, referred to as “effective age” has garnered increasing attention (see e.g., Anderson and Loeser [2], Loeser [24], Li et al. [23], and Spiegelhalter [35]). Effective age represents the average age level of healthy individuals who share the same risk profile as a person being evaluated. In many cases, a person’s health status is more accurately reflected by their effective age rather than their chronological age, as it accounts for the impact of chronic disease risks and overall well-being. Evaluating effective age through gait analysis is a vital component in the proactive management of joint health, facilitating early detection of degeneration, customization of rehabilitation strategies, and the promotion of sustained mobility and independence.
Importantly, coordination among different lower-limb angular movement is pivotal for understanding human movement patterns and their underlying biomechanics. Existing research underscores the importance of inter-joint connectivity in angular kinematics as essential for maintaining balance, stability, and forward progression (see e.g., Ren et al. [32] and Dussault-Picard et al. [10]). Such coordination is crucial for evaluating effective age. As revealed in our analysis, it gives rise to homogeneous age‑associated patterns across different kinematic trajectories, a phenomenon we refer to as kinematic connectivity. Studying homogeneous age‐related associations across kinematic trajectories helps uncover systemic coordinated changes in gait that are more informative and clinically relevant than examining each kinematic trajectory in isolation. It reflects comprehensive aging mechanisms rather than noisy joint‐specific effects. In addition, the connectivity enables simplified model structures and improving evaluation performance.
There are several challenges to solve: 1) the kinematic connectivity structure underlying the association between age and kinematic trajectories is unknown, and should be appropriately identified to achieve accurate and interpretable effective age evaluations. Imposing a fixed connectivity structure as prior knowledge often leads to model misspecification, and such model assumptions are difficult to justify; 2) Since kinematic trajectories are functional data, the kinematic connectivity may change across different dimensions, the developed analytical framework should be tailored to tackle this variability; 3) While kinematic connectivity exists, the idiosyncratic features pertaining to each individual kinematic trajectory also play a crucial role in reflecting lower-limb joint aging. Therefore, distinguishing the shared (homogeneous) components from the trajectory-specific (idiosyncratic) components is essential. All these challenges motivate us to develop a new analytical framework, capable of uncovering kinematic connectivity with complex structures.
To summarize, this study aims to address the following objectives:
  • 1.
    We aim to develop an effective age assessment framework characterizing joint aging mechanisms using kinematic features constructed from angular kinematic trajectories of pelvic, hip, knee, ankle and foot.
  • 2.
    We aim to detect the consistent age‑associated patterns among the angular kinematic trajectories, thereby offering a system‑level view of musculoskeletal aging, and to identify both (partially) homogeneous and idiosyncratic kinematic features that significantly inform effective age evaluation.
@@
Here we analyze a dataset generated from the experiment conducted by Fukuchi et al. [14]. The study involves 42 healthy volunteers aged between 21 and 84, all of whom had no lower-extremity injuries in the six months prior to participation and no orthopedic or neurological conditions affecting their gait. Since the number of participants is limited, it is crucial to avoid building models with high complexity. Studying the association between chronological age and gait angular kinematics in healthy individuals is essential, as it serves as a baseline for assessing new subjects. Although chronological age and effective age do not always align in healthy individuals, particularly in those engaging in regular exercise that can reduce biological age by several years, the phenomenon is not universal and should not be assumed to occur in most healthy people. Data were collected using standard gait-analysis procedures in a 10×12-m room at the Laboratory of Biomechanics and Motor Control at the Federal University of ABC, Brazil. This facility is equipped with a motion-capture system that had 12 cameras, five force platforms embedded in the floor, and a dual-belt instrumented treadmill. Our study focuses on treadmill walking data, as they offer greater consistency and are not influenced by terrain variations. The angular kinematics data were recorded at various gait speeds, categorized into eight levels (S1 — S8) ranging from the slowest to the fastest speeds. The ensemble averages of angular kinematics across different speed levels are shown in Figure 1.
@@
Figure 1: Angular kinematics (right-hand side) during treadmill walking averaged over different levels of gait speed. S5 represents the comfortable walking speed.
@@ 1.2 Functional Linear Regression and Covariate Grouping
Multivariate functional data analysis (FDA) offers a powerful toolbox for handling multiple complex functional objects. Its primary strength lies in addressing the dual challenges of high dimensionality inherent in individual functional elements and the intricate connectivity among them. Recent advancements in this field have enabled researchers to tackle diverse challenges across various scientific domains, including neuroimaging analysis (e.g., Wang et al. [36] and Jiao et al. [18]), public health (e.g., Zhou et al. [41]), and gait analysis (e.g., Cheng et al. [4], Dannenmaier et al. [8], and Cao et al. [3]). These developments underscore the potential of multivariate FDA in addressing complex high-dimensional data. Among its various tools, functional linear regression (see Ramsay and Silverman [30] and Morris [27] for overviews) stands out as a well-established class of linear models. In this article, we focus on the following model with a scalar response $y_{n}\in\mathbb{R}$ and multiple functional covariates $\{X_{jn}(t)\in L^{2}[0,1]\colon j=1,\ldots,p\}$,
@@
As discussed previously, besides evaluating joint effective age, an important task is to account for the connectivity among different covariates (i.e., the angular kinematic trajectories of various joints). Such connectivity leads to covariates being associated with the response variable (i.e., chronological age) in a homogeneous manner, which is crucial to investigate the homogeneous angular kinematic evolution in relation to age. A reasonable strategy to tackle the connectivity is to group different kinematic trajectories based on the similarity of their association with age, and then extract the homogeneous association patterns based on the grouping structure. This first step is known as grouping pursuit (see Shen and Huang [34], Ke et al. [21], Ma and Huang [26], She et al. [33], and Jiao and Chan [17]). However, ordinary grouping pursuit methods face a fundamental limitation: they assume a uniform grouping structure and thus cannot accommodate variability in kinematic connectivity across different dimensions.
@@ 1.3 Contributions and Novelties
Methodologically, we develop a filtration‑based grouping algorithm for model (1-1), termed “filtrated grouping”. This idea is inspired by filtration, a widely used technique in graph analysis and topological data analysis (see, e.g., Edelsbrunner et al. [11], Lee et al. [22], and Hofer et al. [15]), which iteratively constructs a sequence of nested structures and serves as a fundamental building block for multi-resolution analysis. Rather than isolating the effects of individual covariates, our approach captures both shared (homogeneous) and distinct (idiosyncratic) association patterns across multiple levels of resolution. The underlying principle is that as homogeneous components are progressively extracted from different functional covariates, the remaining parts should increasingly exhibit idiosyncratic associations. The covariates are grouped iteratively in each layer of filtration, leading to dimension-specific grouping structures. This new analysis framework not only enhances model interpretability but also improves the accuracy of effective age evaluation. Filtration techniques have rarely been applied to functional data. To the best of our knowledge, the only related work is Jiao et al. [18], which employs filtration to construct a multi-resolution structure for partial common functional principal components (PCfPC), offering a more inclusive and efficient representation of PCfPCs. This paper is the first work to develop filtration techniques for multiple functional regression.
In the application, we develop an innovative tool to study the association between joint angular kinematics and chronological age, serving as a foundation for effective age evaluation and offering new insights into kinematic connectivity. To achieve this, we use multiple functional regression, where age serves as the response variable and angular kinematic curves are treated as functional predictors. By employing our new grouping algorithm, both shared and idiosyncratic kinematic features are effectively captured in a novel hierarchical manner. The variability in the kinematic connectivity structures is effectively addressed, and our analysis highlights the key kinematic features relevant to the assessment of effective age. The findings deepen our understanding of the coordination in age‑related evolution of angular gait kinematics and provide a valuable tool for assessing chronic lower‑limb joint risks.
The rest of the article is organized as follows. We present the filtrated grouping structure and detail the model estimation and feature extraction procedures in Section 2. In Section 4, we apply our methods to analyze the association between chronological age and angular kinematics data. We discuss the results in Section 5. Simulation results and additional algorithms are presented in the supplementary materials.
2 Methodology
@@ 2.1 Preliminaries
Notationally, let $X\in L^{q}_{H}=L_{H}^{q}(\Omega,\mathcal{A},\mathcal{P})$ indicate that, for some $q>0$, a $H$-valued function $X$ satisfies $E\|X(t)\|^{q}<\infty$. In what follows, we assume that there are one scalar response $y_{n}$ representing chronological age and $p$ functional covariates $\{X_{jn}(t)\in L^{2}_{H}\colon j=1,\ldots,p\}$ representing different kinematic trajectories, where $H=L^{2}[0,1]$ is equipped with the inner product $\langle x,y\rangle=\int_{0}^{1}x(t)y(t)dt$ and the norm $\|x\|^{2}=\int_{0}^{1}x^{2}(t)dt$. It is assumed that $\{y_{n},X_{jn}(t)\colon j=1,\ldots,p\}$ are independent across $n=1,\ldots,N$, where $N$ is the sample size. By Mercer’s theorem, we have the following functional principal component decompositions $\mathcal{C}_{j}(\cdot)=\sum_{d\geq 1}\theta_{jd}\langle\nu_{jd},\cdot\rangle\nu_{jd}(t),$ where $\mathcal{C}_{j}(x)(t)=\int C_{j}(t,s)x(s)ds$ for $x\in L^{2}[0,1]$, $\mu_{j}(t)=\mbox{E}X_{jn}(t)$, and $C_{j}(t,s)=\mbox{E}\{(X_{jn}-\mu_{j})(t)(X_{jn}-\mu_{j})(s)\}$. We assume that all the coefficient functions in (1-1) belong to the space spanned by $\{\nu_{jd}(t)\colon d\geq 1,\ j\geq 1\}$.
@@
in which the basis functions $\{\nu_{jd}(t)\colon d\geq 1\}$ and $\{u_{jd}(t)\colon d\geq 1\}$ can either be orthonormal or not. The infinite matrix $\{M_{j,dd^{\prime}}\colon d,d^{\prime}\geq 1\}$ satisfies the condition $\sum_{d^{\prime}\geq 1}M_{j,d_{1}d^{\prime}}(\int\nu_{jd_{2}}u_{jd^{\prime}})=\bm{1}_{\{d_{1}=d_{2}\}}$. Then model (1-1) can be equivalently rewritten in the following form
@@
For illustrating the new filtrated grouping structure, we focus on the equivalent form (2-2), as the representation is dimension-specific (layer-specific), enabling a more precise characterization of grouping structures across different dimensions.
@@ 2.2 Filtrated Grouping Structure
In this study, we model kinematic connectivity using a hierarchical grouping strategy. To begin, we introduce the ordinary grouped regression model as a foundation. Given a grouping structure $G=(\mathcal{K}_{1},\ldots,\mathcal{K}_{m})$ satisfying $\mathcal{K}_{i}\cap\mathcal{K}_{i^{\prime}}=\emptyset$ for $i\neq i^{\prime}$ and $\cup_{i=1}^{m}\mathcal{K}_{i}=\{1,\ldots,p\}$, the ordinary grouped model based on coefficient homogeneity is given as follows
@@
where $Z_{in}(t)=\sum_{j\in\mathcal{K}_{i}}X_{jn}(t)$ and $\alpha_{i}(t)$ is the common coefficient function of group $\mathcal{K}_{i}$, capturing the homogeneous association shared by the covariates within the group. The ordinary model (1-1) cannot account for the homogeneity of functional coefficients and the estimation burden is high even for moderate $p$. Although the grouped model (2-3) alleviates the estimation burden and improve model interpretability, it fails to explain the data mechanism when the grouping structures are different across different dimensions. Indeed, as demonstrated in Section 4, both common and idiosyncratic components of different angular kinematic trajectories play important roles in effective age evaluation, and imposing a uniform grouping structure results in suboptimal model performance. This motivates the development of the filtrated grouping procedure.
Building upon the equivalent model representation (2-2), we develop the following filtrated grouped model
@@
where $\mathcal{K}_{d,i}$ denotes the $i$-th group in the $d$-th dimension/filtration layer and $m_{d}$ is the number of groups in the $d$-th dimension. Compared to the ordinary grouped model (2-3), the filtrated grouped model (2-4) accommodates grouping variability across different dimensions. The homogeneous components are defined as follows
$\zeta_{nd}^{(i)}\overset{\vartriangle}{=}\sum_{j\in\mathcal{K}_{d,i}}\xi_{jn,d},\ i\geq 1,\ d\geq 1.$
These components capture the features shared among covariates within each group, which are expected to have strong associations with the response. In particular, when $\mathcal{K}_{d,i}$ contains only a single covariate, $\zeta_{nd}^{(i)}$ captures the idiosyncratic feature of the covariate. In addition to obtaining these homogeneous components, we also need to estimate the associated coefficient scores $\{a_{d,i}\colon i\geq 1,\ d\geq 1\}$.
The homogeneous components are extracted in a hierarchical manner, resulting in group indices $\{\mathcal{K}_{d,i}\colon i\geq 1,d\geq 1\}$ that follow a forest structure (see Figure 2). In the first layer, we identify the most homogeneous covariate components and group them accordingly to obtain $\{\zeta_{n1}^{(i)}\colon i\geq 1\}$ and estimate $\{a_{1,i}\colon i\geq 1\}$. After filtering out these homogeneous components from each covariate, we divide $\{\mathcal{K}_{1,i}\colon i\geq 1\}$ into $\{\mathcal{K}_{2,i}\colon i\geq 1\}$, which are either nested within or identical to the groups in the previous layer (i.e., for any $i\geq 1$, $\mathcal{K}_{2,i}\subseteq\mathcal{K}_{1,i^{\prime}}$ for some $i^{\prime}\geq 1$), and then obtain the homogeneous components in the new layer (i.e., $\{\zeta_{n2}^{(i)}\colon i\geq 1\}$) and estimate the associated coefficient scores (i.e., $\{a_{2,i}\colon i\geq 1\}$). The splitting procedure continues until a predefined stopping criterion is met. In the deeper layers of the filtration, higher-resolution homogeneous components, constructed from fewer covariates, are progressively extracted.
Figure 2: Hierarchical forest structure of coefficient homogeneity. Here, $L_{d}$ signifies the $d$-th layer of filtration. Different groups are represented by different colored solid circles. Each group is either nested within or identical to its “parent” group (i.e., the group from which it originates through a split) in the previous layer.
In the remainder of the section, we address the following problems: Given the forest structure, how to extract the homogeneous components and estimate the coefficient scores? How to determine a reasonable forest structure for filtrated covariate grouping? In Section 2.3, we address the first problem by developing a filtrated functional partial least squares approach. In Section 2.4, we construct the forest structure based on the grouping path derived from the grouping pursuit problem.